# BELT: Old-School Backdoor Attacks can Evade the State-of-the-Art Defense with Backdoor Exclusivity Lifting

### v1

@@ Abstract
In this paper, we propose and investigate a new characteristic of backdoor attacks, namely, backdoor exclusivity, which measures the ability of backdoor triggers to remain effective in the presence of input variation. Building upon the concept of backdoor exclusivity, we propose Backdoor Exclusivity LifTing (BELT), a novel technique which suppresses the association between the backdoor and fuzzy triggers to enhance backdoor exclusivity for defense evasion. Extensive evaluation on three popular backdoor benchmarks validate, our approach substantially enhances the stealthiness of four old-school backdoor attacks, which, after backdoor exclusivity lifting, is able to evade six state-of-the-art backdoor countermeasures, at almost no cost of the attack success rate and normal utility. For example, one of the earliest backdoor attacks BadNet, enhanced by BELT, evades most of the state-of-the-art defenses including ABS and MOTH which would otherwise recognize the backdoored model.
@@ 1 Introduction
However, backdoor robustness also results in a numerous number of approximate triggers that can unintentionally activate the backdoor. This increases the possibility of the backdoor to be detected by a model inspector. Most of the state-of-the-art backdoor defenses [16, 33, 51, 11, 58, 49] do transform the robustness advantage of the backdoor into a weakness for trigger proximity. A prerequisite for the successful implementation of these backdoor defenses is the ability to effectively find an approximate replacement for the original trigger, leveraging methods such as reverse engineering [51], sample overlay [16], or attribution techniques [11]. For instance, MOTH published at S&P 2022 [49] achieves model orthogonalization by relearning reverse triggers and natural triggers, which results effective backdoor elimination. Section B provides more detailed analysis. As shown in Figure 1, despite the triggers identified by these defense strategies not being identical to the original trigger, they still possess the ability to activate the backdoor and can thus be classified as fuzzy triggers.
@@
$\bullet$ On the basis of backdoor exclusivity, we propose a new backdoor attack technique called BELT, which suppresses the association between the backdoor and fuzzy triggers to enhance backdoor exclusivity for defense evasion. Our technique can be seamless combined with existing backdoor attacks and enhance the classical attacks to evade the SOTA backdoor defenses which would otherwise detect and eliminate them.
$\bullet$ We comprehensively evaluate the performance of BELT on six SOTA backdoor defenses, and the results validate that BELT helps four classical backdoor attacks to evade the SOTA backdoor defenses, while preserving the attack success rate and the clean accuracy. This confirms that highly exclusive backdoors are more difficult to be detected. To facilitate future studies, we open-source our code in the following repository: https://anonymous.4open.science/r/BELT-7492/.
@@ 3 Problem Definition
Specifically, for any perturbation $\delta\in\mathbb{R}^{d}$ with $\|\delta\|_{p}>\|\delta_{max}\|_{p}$, it holds that if $c(F_{\theta^{\prime}}(x))\neq y_{t}$ then $c(F_{\theta^{\prime}}(x_{t^{\prime}}))\neq c(F_{\theta^{\prime}}(x_{t}))$. With a trigger upper bound $\|\delta_{max}\|_{p}$ for a poisoned input $x_{t}$, we can devise a metric Excl to measure its exclusivity, which can be expressed as:
$\text{Excl}=1-\frac{\|\delta_{max}\|_{p}}{\|\delta_{\text{b}}\|_{p}}.$ (3)
In practice, we default to using the Euclidean norm ($l_{2}$-norm) for exclusivity. The exclusivity metric $\text{Excl}\in[0,1]$ is used to quantify the uniqueness of triggers. As shown in  Figure 2, the origin represents the original trigger, the coordinate axis represents the size of the disturbance suffered by the original trigger, the black circle is the disturbance boundary $\delta_{b}$, and the red and blue points represent the fuzzy trigger and ineffective trigger the respectively. For a backdoor with exclusivity approaching 0%, $\beta_{U}$ approaches the perturbation boundary $\delta_{b}$. In this case, there is no clear trigger upper bound $\beta_{U}$ to ensure trigger ineffectiveness, implying that all perturbed triggers may be fuzzy. For a backdoor with exclusivity approaching 100%, $\beta_{U}$ approaches 0. This can be interpreted as all perturbed triggers with $\|\delta\|_{p}>0$ are ineffective, indicating that the backdoor only allows activation by the original trigger, making the trigger unique.
@@ 4 Security Settings
Attack Scenarios and Capabilities. We consider two common scenarios for backdoor attacks: data outsourcing [9, 11] and model outsourcing [18, 33]. In the data outsourcing scenario, the victim can train the neural network model but faces challenges due to insufficient training data. Consequently, the victim outsources data to a third party (e.g., Amazon Mechanical Turk [50]) or collects data from public internet resources [26]. In this scenario, the attacker can provide toxic data directly to the victim or spread poisoned data over the internet. However, the attacker cannot control the training process. Despite having access to the entire training data, the victim can implement countermeasures like data cleaning to filter out suspicious data [16, 11].
In a model outsourcing scenario, the victim may lack domain expertise and/or computational resources but has training resources. As a result, the victim may consider outsourcing model training to a third party that is not fully trusted [26] or directly using third-party models provided by popular platforms such as HuggingFace [20]. In this scenario, the attacker can manipulate the data and control the entire model training process. However, the victim does not have access to the tampered training data but can use a small portion of the benign data to implement countermeasures such as backdoor detection or elimination to mitigate the risks associated with outsourced models [35, 49].
@@ 5.1 Methodology Overview
Pipeline Overview. The entire injection process of BELT is illustrated in  Figure 4. Similar to previous data poisoning-based backdoor attacks [9, 18, 25, 39], we initially generate triggers and modify labels to $y_{t}$ to create a set of poisoned samples $D_{t}$. To improve the backdoor exclusivity, we randomly mask the triggers to produce fuzzy triggers, forming a set of cover samples $D_{c}$. It is important to note that, to sever the link between fuzzy triggers and the backdoor, the labels of cover samples remain unchanged. Subsequently, the poisoned, cover, and benign samples from the original training set are employed as training data for the backdoored model. We consider two prevalent attack scenarios for BELT: data outsourcing and model outsourcing. In the data outsourcing scenario, attackers can influence the training data but cannot manipulate the loss function. The backdoored model in the data outsourcing scenario is optimized using the loss formula Equation 8. In the model outsourcing scenario, attackers can further enhance the backdoor attack effect by utilizing the momentum center loss Equation 10 to increase the distance between triggers and blurred triggers in the feature space. Finally, the backdoored model is released into public communities such as HuggingFace or GitHub, leading to unpredictable consequences [20].
@@ 5.2 Backdoor Exclusivity
Optimization direction. We first calculate the maximum perturbation of the trigger within a valid numerical range, referred to as the perturbation boundary $\delta_{b}$ (as defined in Definition 1). We use this perturbation boundary as the guiding direction for the perturbation $\delta$ during the optimization process. This enables the perturbation to be adjusted along the direction of the perturbation boundary, providing more optimization space to find a trigger with a larger perturbation.
@@
$\text{Excl}=\frac{1}{n}\sum_{i=1}^{n}(1-\frac{\|\delta_{i,max}\|_{2}}{\|\delta_{i,b}\|_{2}})$ (7)
where $\|\delta_{i}\|$ and $\|\delta_{i,b}\|$ represent the trigger upper bound and perturbation boundary for the $i$-th sample, respectively. The exclusivity average of the $n$ poisoned samples is used to represent the exclusivity of the backdoor.
@@ 6.1 Overview of Evaluation
We evaluate the effectiveness of the BELT on three datasets (refer to Table I) using four representative backdoor attacks (as detailed in Section 6.2). These attacks range from input-agnostic to input-aware. We also assess the proposed exclusivity metric for these backdoor attacks. Furthermore, we enhance the efficacy of BELT under six SOTA backdoor defense frameworks (discussed in Section 6.3).
@@
Training Setup. For model training, we configure the batch size to 128 and employ the SGD optimizer with a learning rate of $1\times 10^{-1}$ for 100 epochs of optimization. In the case of input-aware backdoor attacks, namely ISSBA and IAD, we adhere to the encoder structures specified in their respective original papers. In the assessment of backdoor defenses, we follow the settings outlined in the original papers. For optimizing the exclusivity, we utilize the Adam optimizer with a learning rate set to $1\times 10^{-3}$. Specifically, we set the poison rate for BadNet and Blended to 0.01. For IAD and ISSBA, the poison rates are set to 0.02 and 0.1, respectively. Cover samples are constructed based on the dirty samples, and the mask rate of triggers is set to 0.2.
@@
BadNet[18] 93.17 100.00 0.00 97.12 99.15 0.00 50.95 100.00 1.98
${}^{+}$ 94.85 99.98 16.45 98.12 99.99 16.24 52.79 99.99 5.22
${}^{++}$ 94.62 100.00 17.88 98.35 99.98 19.38 52.31 100.00 5.99
Blended[9] 95.28 99.79 0.09 98.37 99.27 0.36 51.77 99.58 1.55
${}^{+}$ 94.48 98.47 1.58 96.41 99.37 2.07 51.95 98.78 3.01
${}^{++}$ 94.61 99.14 2.00 97.70 99.98 2.47 51.31 99.29 3.44
ISSBA[27] 94.80 99.03 36.10 97.64 99.89 38.10 48.70 100.00 39.78
${}^{+}$ 95.16 97.57 38.66 98.12 99.97 41.57 48.86 99.91 43.28
${}^{++}$ 94.13 99.97 43.79 98.01 100.00 47.02 48.26 99.87 49.42
IAD[39] 93.32 99.93 0.53 97.19 99.25 1.33 50.43 99.18 3.93
${}^{+}$ 92.78 97.04 2.89 97.33 99.51 5.44 47.95 99.47 6.47
${}^{++}$ 93.34 99.56 3.74 97.74 99.85 10.93 48.93 98.95 16.97
@@ 6.2 BELT Evaluation and Analysis
We evaluate the BELT framework in two scenarios, i.e., data outsourcing and model outsourcing, by combining BELT with four SOTA attacks. In the data outsourcing scenario, we introduce dirty and cover samples through data poisoning alone, denoted as ${}^{+}$. In the model outsourcing scenario, where the attacker has complete control over the entire model training process, we enhance the BELT with momentum-centered loss (Equation 11), denoted as ${}^{++}$. The summarized experimental results are presented in Table II.
@@
BadNet. To evaluate BadNet[18], we employ $6\times 6$ random blocks of pixels as triggers, as depicted in the first column of Figure 3. Utilizing BELT, we strengthen BadNet’s exclusivity under data outsourcing and attack outsourcing scenarios, denoted as ${}^{+}$ and ${}^{++}$, respectively. From Table II, it is evident that both BadNet before and after reinforcement exhibit high ASR while maintaining a comparable CDA to the clean model. However, analyzing the backdoor exclusivity reveals a significant distinction. The regular BadNet’s Excl is notably low, approaching 0%, indicating loose activation conditions and numerous fuzzy triggers that may activate the backdoor. In contrast, the exclusivity of ${}^{+}$ and ${}^{++}$ shows substantial improvement compared to regular BadNet, increasing from 0% to 5%-20%. This enhancement implies that ${}^{+}$ and ${}^{++}$ respond more accurately to triggers, reducing the occurrence of potentially fuzzy triggers. We elucidate the reasons behind the regular BadNet’s ease of detection from the exclusivity perspective and demonstrate that exclusivity-enhanced BadNet successfully evades SOTA backdoor defenses, as detailed in Section 6.3.
@@
Blended. To evaluate Blended[9], we employ the hello_kitty image with transparency to 0.2 as triggers, as depicted in the second column of Figure 3. Similar to BadNet evaluation, we conduct experiments under data outsourcing and model outsourcing scenarios, denoted as ${}^{+}$ and ${}^{++}$. As shown in Table II, BELT demonstrates negligible effects on CDA and ASR, while significantly improving the backdoor exclusivity, increasing from 0.09% to 2%-3%. It is noteworthy that the exclusivity of Blended is relatively smaller than BadNet, attributed to the broader range of perturbations and fixed perturbation value of Blended trigger, the trigger itself has lower exclusivity compared to BadNet, resulting in numerous potential fuzzy triggers. However, BELT still significantly increases the exclusivity of Blended by approximately 3-20 times and manifests in corresponding evasion effects in the MOTH defense (Section 6.3).
@@
ISSBA. ISSBA aims to propose an imperceptible backdoor trigger[25]. We employ a DNN-based image steganography network to generate sample-specific invisible additive noises as backdoor triggers by encoding an attacker-specified string into benign images, as shown in the third column of Figure 3. According to our experimental results, ${}^{+}$ and ${}^{++}$ preserve ASR and have minimal impact on CDA. The exclusivity of ${}^{+}$ and ${}^{++}$ still outweighs ISSBA. It is noted that the Excls of ISSBA are significantly greater than other attacks. This phenomenon can be attributed to two folds. Firstly, ISSBA adds very small perturbation to the images, (about -0.05 to 0.05), making them imperceptible to human eyes. This significantly reduces the range for perturbing the trigger, resulting in fewer potential fuzzy triggers and inherently higher exclusivity compared to traditional visible perturbations. Secondly, ISSBA is a sample-specific backdoor attack. Intuitively, sample-specific attacks create sample-specific triggers, which form a corresponding input-dependent trigger space. This further limits the range in which the trigger can be perturbed, thereby enhancing exclusivity.
@@
IAD. To evaluate IAD[40], we train a 7-layer CNN-based generator for three datasets from scratch to generate input-aware triggers for dirty samples. As documented in Table II, CDA and ASR of ${}^{+}$ and ${}^{++}$ remain largely unchanged compared to IAD. The exclusivity of ${}^{+}$ and ${}^{++}$ show improvement compared to IAD. In contrast to ISSBA, IAD adds more pronounced perturbations to images. The lower exclusivity indicates a wider range of perturbation endurance. However, as analyzed in ISSBA, input-aware triggers possess higher exclusivity naturally compared to Blended. Therefore, ${}^{+}$ and ${}^{++}$ possess Excls between 0.53 and 16.97.

### v2

@@ Abstract
In this paper, we propose and investigate a new characteristic of backdoor attacks, namely, backdoor exclusivity, which measures the ability of backdoor triggers to remain effective in the presence of input variation. Building upon the concept of backdoor exclusivity, we propose Backdoor Exclusivity LifTing (BELT), a novel technique which suppresses the association between the backdoor and fuzzy triggers to enhance backdoor exclusivity for defense evasion. Extensive evaluation on three popular backdoor benchmarks validate, our approach substantially enhances the stealthiness of four old-school backdoor attacks, which, after backdoor exclusivity lifting, is able to evade seven state-of-the-art backdoor countermeasures, at almost no cost of the attack success rate and normal utility. For example, one of the earliest backdoor attacks BadNet, enhanced by BELT, evades most of the state-of-the-art defenses including ABS and MOTH which would otherwise recognize the backdoored model.
@@ 1 Introduction
However, backdoor robustness also results in a numerous number of approximate triggers that can unintentionally activate the backdoor. This increases the possibility of the backdoor to be detected by a model inspector. Most of the state-of-the-art backdoor defenses [17, 35, 52, 12, 59, 50] do transform the robustness advantage of the backdoor into a weakness for trigger proximity. A prerequisite for the successful implementation of these backdoor defenses is the ability to effectively find an approximate replacement for the original trigger, leveraging methods such as reverse engineering [52], sample overlay [17], or attribution techniques [12]. For instance, MOTH published at S&P 2022 [50] achieves model orthogonalization by relearning reverse triggers and natural triggers, which results in effective backdoor elimination. Section B provides more detailed analysis. As shown in Figure 1, despite the triggers identified by these defense strategies not being identical to the original trigger, they still possess the ability to activate the backdoor and can thus be classified as fuzzy triggers.
@@
$\bullet$ On the basis of backdoor exclusivity, we propose a new backdoor attack technique called BELT, which suppresses the association between the backdoor and fuzzy triggers to enhance backdoor exclusivity for defense evasion. Our technique can be seamlessly combined with existing backdoor attacks and enhance the classical attacks to evade the SOTA backdoor defenses which would otherwise detect and eliminate them.
$\bullet$ We comprehensively evaluate the performance of BELT on seven SOTA backdoor defenses, and the results validate that BELT helps four classical backdoor attacks to evade the SOTA backdoor defenses, while preserving the attack success rate and the clean accuracy. This confirms that highly exclusive backdoors are more difficult to be detected. To facilitate future studies, we open-source our code in the following repository: https://github.com/JSun20220909/BELT.
@@ 3 Problem Definition
Specifically, for any perturbation $\delta\in\mathbb{R}^{d}$ with $\|\delta\|_{p}>\|\delta_{max}\|_{p}$, it holds that if $c(F_{\theta^{\prime}}(x))\neq y_{t}$ then $c(F_{\theta^{\prime}}(x_{t^{\prime}}))\neq c(F_{\theta^{\prime}}(x_{t}))$. With a trigger upper bound $\|\delta_{max}\|_{p}$ for a poisoned input $x_{t}$, we can devise a metric Excl to measure its exclusivity. Considering that different triggers have different number of pixels $N$ and range of values [$t_{min}$, $t_{max}$], exclusivity can be expressed as:
${\color[rgb]{0,0,0}\text{Excl}=1-(\frac{\|\delta_{max}\|_{p}}{\|\delta_{b}\|_{p}})^{(t_{max}-t_{min})\sqrt{N}}.}$ (3)
In practice, we default to using the Euclidean norm ($l_{2}$-norm) for exclusivity. The exclusivity metric $\text{Excl}\in[0,1]$ is used to quantify the uniqueness of triggers. As shown in  Figure 2, the origin represents the original trigger, the coordinate axis represents the size of the disturbance suffered by the original trigger, the black circle is the disturbance boundary $\delta_{b}$, and the red and blue points respectively represent the fuzzy trigger and ineffective trigger. For a backdoor with exclusivity approaching 0%, $\beta_{U}$ approaches the perturbation boundary $\delta_{b}$. In this case, there is no clear trigger upper bound $\beta_{U}$ to ensure trigger ineffectiveness, implying that all perturbed triggers may be fuzzy. For a backdoor with exclusivity approaching 100%, $\beta_{U}$ approaches 0. This can be interpreted as all perturbed triggers with $\|\delta\|_{p}>0$ are ineffective, indicating that the backdoor only allows activation by the original trigger, making the trigger unique.
@@ 4 Security Settings
Attack Scenarios and Capabilities. We consider two common scenarios for backdoor attacks: model outsourcing [10, 12] and data outsourcing [19, 35]. In the model outsourcing scenario, the victim may lack domain expertise and/or computational resources but has data resources. As a result, the victim may consider outsource the model training to a third party which may not be fully trusted [27] or directly use third-party models provided by popular platforms such as HuggingFace [21]. In this scenario, the attacker enhances backdoors by manipulating the data and adding loss terms. For defenses, the victim does not have access to the tampered training data, but does have access to the trained model. The victim may consider implement model-level defenses based on a small portion of benign data, such as Fine-Pruning [34], NC [52], ABS [35], MNTD [59], and MOTH [50], to mitigate the risks associated with outsourcing models.
In the data outsourcing scenario, the victim has computational resources and expertise to train the model but faces challenges due to insufficient training data. Consequently, the victim may consider outsource the training data preparation to a third party (e.g., Amazon Mechanical Turk [51]) or collect data from public internet resources [27]. In this scenario, the attacker would provide toxic data directly to the victim or spread poisoned data over the internet, but has no capability to manipulate the training process of the model. Therefore, for defenses, the victim has access to the training data and the model, but has difficulty in ensuring the dataset is clean. Therefore, the victim may consider data-level defenses to filter the poisoned samples, such as STRIP [17] and SentiNet [12], and employ model-level defenses only when a small clean dataset is available.
@@ 5.1 Methodology Overview
Pipeline Overview. The entire injection process of BELT is illustrated in  Figure 4. Regular backdoor attacks [10, 19, 26, 40] involve mixing poison samples ($D_{t}$) which carry triggers (e.g., a white circle) and incorrect labels into the training set. Beyond this, BELT also mixes cover samples ($D_{c}$) with fuzzy triggers (e.g., incomplete white circles) and correct labels to enhance the backdoor exclusivity. The training set comprises the clean dataset $D$, the poison dataset $D_{t}$, and the cover dataset $D_{c}$, which implant a highly exclusive backdoor in the victim model through standard training procedures. We consider two prevalent attack scenarios for BELT: data outsourcing and model outsourcing. In the data outsourcing scenario, attackers can influence the training data but cannot manipulate the loss function. The backdoored model in the data outsourcing scenario is optimized using the loss formula Equation 8. In the model outsourcing scenario, attackers can further enhance the backdoor attack effect by utilizing the momentum center loss Equation 10 to increase the distance between triggers and blurred triggers in the feature space. Finally, the backdoored model is released into public communities such as HuggingFace or GitHub, leading to unpredictable consequences [21].
@@ 5.2 Backdoor Exclusivity
Optimization direction. We first calculate the maximum perturbation of the trigger within a valid numerical range, referred to as the perturbation boundary $\delta_{b}$ (as defined in Definition 1). It is worth noting that the valid numerical range is the permissible pixel value of the trigger, denoted as [$t_{min}$, $t_{max}$], which varies according to different backdoor definitions. We use this perturbation boundary as the guiding direction for the perturbation $\delta$ during the optimization process. This enables the perturbation to be adjusted along the direction of the perturbation boundary, providing more optimization space to find a trigger with a larger perturbation.
@@
${\color[rgb]{0,0,0}\text{Excl}=\frac{1}{n}\sum_{i=1}^{n}(1-(\frac{\|\delta_{i,max}\|_{2}}{\|\delta_{i,b}\|_{2}})^{(t_{i,max}-t_{i,min})\sqrt{N_{i}}}),}$ (7)
where $\|\delta_{i,max}\|$ and $\|\delta_{i,b}\|$ represent the trigger upper bound and perturbation boundary for the $i$-th sample, respectively. The exclusivity average of the $n$ poisoned samples is used to represent the exclusivity of the backdoor.
@@ 6.1 Overview of Evaluation
We evaluate the effectiveness of the BELT on three datasets (refer to Table I) using four representative backdoor attacks (as detailed in Section 6.2). These attacks range from input-agnostic to input-aware. We also assess the proposed exclusivity metric for these backdoor attacks. Furthermore, we enhance the efficacy of BELT under seven SOTA backdoor defense frameworks (discussed in Section 6.3).
@@
Training Setup. For model training, we configure the batch size to 128 and employ the SGD optimizer with a learning rate of $1\times 10^{-1}$ for 100 epochs of optimization. In the case of input-aware backdoor attacks, namely ISSBA and IAD, we adhere to the encoder structures specified in their respective original papers. In the assessment of backdoor defenses, we follow the settings outlined in the original papers. For optimizing the exclusivity, we utilize the Adam optimizer with a learning rate set to $1\times 10^{-3}$. Specifically, we set the poison rate for BadNet and Blended to 0.01. For IAD and ISSBA, the poison rates are set to 0.02 and 0.1, respectively. Cover samples are constructed based on the cover triggers, and the mask rate of triggers is set to 0.2.
@@
BadNet[19] 93.17 100.00 0.00 97.12 99.15 0.00 50.95 100.00 34.01
BadNet+ 94.85 99.98 84.55 98.12 99.99 84.14 52.79 99.99 67.19
BadNet++ 94.62 100.00 87.09 98.35 99.98 89.34 52.31 100.00 72.30
Blended[10] 95.28 99.79 0.99 98.37 99.27 3.92 51.77 99.58 29.27
Blended+ 94.48 98.47 16.18 96.41 99.37 20.70 51.95 98.78 49.22
Blended++ 94.61 99.14 20.06 97.70 99.98 24.21 51.31 99.29 53.98
ISSBA[28] 94.80 99.03 21.98 97.64 99.89 23.34 48.70 100.00 43.00
ISSBA+ 95.16 97.57 23.73 98.12 99.97 25.76 48.86 99.91 46.66
ISSBA++ 94.13 99.97 27.33 98.01 100.00 29.68 48.26 99.87 53.03
IAD[40] 93.32 99.93 6.18 97.19 99.25 14.84 50.43 99.18 38.19
IAD+ 92.78 97.04 29.67 97.33 99.51 48.89 47.95 99.47 55.19
IAD++ 93.34 99.56 36.71 97.74 99.85 75.07 48.93 98.95 89.26
@@ 6.2 BELT Evaluation and Analysis
We evaluate the BELT framework in two scenarios, i.e., data outsourcing and model outsourcing, by combining BELT with four SOTA attacks. In the data outsourcing scenario, we introduce dirty and cover samples through data poisoning alone, denoted as +. In the model outsourcing scenario, where the attacker has complete control over the entire model training process, we enhance the BELT with momentum-centered loss (Equation 11), denoted as ++. The summarized experimental results are presented in Table II.
@@
BadNet. To evaluate BadNet[19], we employ $6\times 6$ random blocks of pixels as triggers, as depicted in the first column of Figure 3. Utilizing BELT, we strengthen BadNet’s exclusivity under data outsourcing and attack outsourcing scenarios, denoted as BadNet+ and BadNet++, respectively. From Table II, it is evident that both BadNet before and after reinforcement exhibit high ASR while maintaining a comparable CDA to the clean model. However, analyzing the backdoor exclusivity reveals a significant distinction. The regular BadNet’s Excl is notably low, approaching 0%, indicating loose activation conditions and numerous fuzzy triggers that may activate the backdoor. In contrast, the exclusivity of BadNet+ and BadNet++ shows substantial improvement compared to regular BadNet, increasing from 0% to 84%-87%. This enhancement implies that BadNet+ and BadNet++ respond more accurately to triggers, reducing the occurrence of potentially fuzzy triggers. We elucidate the reasons behind the regular BadNet’s ease of detection from the exclusivity perspective and demonstrate that exclusivity-enhanced BadNet successfully evades SOTA backdoor defenses, as detailed in Section 6.3.
@@
Blended. To evaluate Blended[10], we employ the hello_kitty image with transparency to 0.2 as triggers, as depicted in the second column of Figure 3. Similar to BadNet evaluation, we conduct experiments under data outsourcing and model outsourcing scenarios, denoted as Blended+ and Blended++. As shown in Table II, BELT demonstrates negligible effects on CDA and ASR, while significantly improving the backdoor exclusivity, increasing from 0.99% to 16%-20%. It is noteworthy that the exclusivity of Blended is relatively smaller than BadNet, attributed to the broader range of perturbations and fixed perturbation value of Blended trigger, the trigger itself has lower exclusivity compared to BadNet, resulting in numerous potential fuzzy triggers. However, BELT still significantly increases the exclusivity of Blended by approximately 2-20 times and manifests in corresponding evasion effects in the MOTH defense (Section 6.3).
@@
ISSBA. ISSBA aims to propose an imperceptible backdoor trigger[26]. We employ a DNN-based image steganography network to generate sample-specific invisible additive noises as backdoor triggers by encoding an attacker-specified string into benign images, as shown in the third column of Figure 3. According to our experimental results, ISSBA+ and ISSBA++ preserve ASR and have minimal impact on CDA. The exclusivity of ISSBA+ and ISSBA++ still outweighs ISSBA. It is noted that the Excls of regular ISSBA are significantly greater than other attacks. This phenomenon can be attributed to two folds. Firstly, ISSBA adds very small perturbation to the images, (about -0.05 to 0.05), making them imperceptible to human eyes. This significantly reduces the range for perturbing the trigger, resulting in fewer potential fuzzy triggers and inherently higher exclusivity compared to traditional visible perturbations. Secondly, ISSBA is a sample-specific backdoor attack. Intuitively, sample-specific attacks create sample-specific triggers, which form a corresponding input-dependent trigger space. This further limits the range in which the trigger can be perturbed, thereby enhancing exclusivity.
@@
IAD. To evaluate IAD[41], we train a 7-layer CNN-based generator for three datasets from scratch to generate input-aware triggers for dirty samples. As documented in Table II, CDA and ASR of IAD+ and IAD++ remain largely unchanged compared to IAD. The exclusivity of IAD+ and IAD++ show improvement compared to IAD. In contrast to ISSBA, IAD adds more pronounced perturbations to images. However, as analyzed in ISSBA, input-aware triggers possess higher exclusivity naturally compared to Blended. Therefore, IAD+ and IAD++ possess Excls between 29% and 89%.